OpenStudy (anonymous):

two flasks are connected with a stopwingspan. flask #1 has a volume of 2.5L and contains oxygen gas at a pressure of .7atm. Flask #2 has a colume of 3.8L and contains hydrogen gas at a pressure of 1.25atm. when the stopwingspan between the two flasks is opened and the gases mix, what will the resulting pressure of the gas mixture be?

OpenStudy (zpupster):

how you do this is add the 2 volumes 2.5+3.8=6.3 and call this V2 and then use the formula P1V1=P2V2 for flask 1 P1 = .7atm V1 = 2.5L Pf1 = ? V2 = .63L rearranging the formula Pf1= (.7)(2.5)/6.3 = ? the same for flask2 P1 = 1.25atm V1 = 3.8L Pf2 = ? V2 = 6.3L Pf2=(1.25)(3.8)/6.3 lastly Pf1 + Pf2 = Ptot Ptot -= ?

OpenStudy (beccaboo333):

I feel like this is more of a math question but at the same time im not sure. @zpupster explained this well.